OpenStudy (anonymous):

A pendulum is released and swings until it stops. If it passes through an arc of 35 inches the first pass, and if on each successive pass it travels 4/5 the distance of the preceding pass, how far will it travel before stopping?

OpenStudy (anonymous):

possible answers are, 210,175,140, and 315

OpenStudy (anonymous):

I think there isn't an answer since if it always travels 80% of the previous distance, then technically it should always be travelling

OpenStudy (anonymous):

I tired taking the log of it but ended up having to take log 0, which is undefined

OpenStudy (psymon):

You wouldn't be doing anything close to that, it's a geometric series problem.

OpenStudy (anonymous):

Aaa right I misread the question

OpenStudy (anonymous):

Yeah you're right it should converge

OpenStudy (anonymous):

that's right it's a geometric series problem, I'm just trying to figure out how to set it up

OpenStudy (psymon):

I'm just making sure I don't put down something stupid as an answer. Bit rusty.

OpenStudy (anonymous):

It's multiple choice so I was going to just pick one.

OpenStudy (anonymous):

Well the basic formula for a geometric series is \[Sn = \frac{ a1(1-r^n) }{ 1-r }\]

OpenStudy (anonymous):

So all you need to do is insert the values into their places

OpenStudy (psymon):

So you start with 35 and then each pass is 35(4/5)^n. So you could have 35 + 35*sum (4/5)^n. You can use your geometric series formula of 1/(1-r) to see what (4/5)^n converges to from there. If you do that, 1/(1-(4/5)) = 5. So you end up 35 + 35(5) = 210

OpenStudy (anonymous):

\[Sn = \frac{ 35(1-0.8^n) }{ 1-0.8 }\]

OpenStudy (anonymous):

And then you have to solve for n

OpenStudy (psymon):

Here's how I set it up personally, maybe it'll look like something familiar, maybe not *draws*

OpenStudy (anonymous):

Teinis that would give me 35

OpenStudy (anonymous):

Nah I don't think it would

OpenStudy (anonymous):

Because you can't just cancel 0ff the (1-0.8)

OpenStudy (psymon):